### v2

# Characterization of symmetries of contact Hamiltonian systems

Abstract
In this paper, we explore the relationship between dynamical symmetries, Cartan symmetries, and dynamical similarities in contact mechanics. Using an alternative decomposition of vector fields, we provide a characterization of those symmetries and a new description in terms of tensor densities. Additionally, we show that, in certain cases, this construction enables the recovery of integrals of motion.
Keywords: contact geometry, dynamical symmetries, Cartan symmetries, scaling symmetries, dynamical similarities, Hamiltonian systems
MSC2010: 70G45, 70G65, 70H33, 34A26, 53D10
1 Introduction
Contact Hamiltonian systems are a natural generalization of the usual symplectic Hamiltonian systems to odd-dimensional manifolds [2005, Arnold2009]. The move to odd-dimensional spaces inevitably leads to the loss of some properties and the emergence of new dynamical phenomena. For example, it is possible to use contact hamiltonians to describe dynamical system with linear dissipation in a rather immediate way [Bravetti2017]. This has attracted the attention of the community, as it would allow to describe a wider range of physical systems, leading to multiple efforts towards the development of a comprehensive contact geometric description of dissipative dynamical systems [Bravetti2015, Bravetti2017, Leon2019].
Despite the differences that distinguish genuinely Hamiltonian systems from their contact counterparts, with some care it is possible to extend surprisingly many results across the two settings. For example, it is possible to generalize Noether theorem [Bravetti2021, Bravetti2023a], introduce analogues to conserved quantities [Leon2019] and even define a notion of integrability.
For contact systems any $1$-parameter transformation group that preserves the contact structure is a contact Hamiltonian action [geiges]. However, in this setting the conservation of the distribution does not imply a conservation of quantities along the motion, and the definition and characterization of symmetries and invariants becomes much more involved. Indeed, there is no unique natural way to define the notion of symmetry of a contact Hamiltonian system, and in the past researchers have already explored different definitions and their properties [Bravetti2021, Bravetti2023a, Bravetti2017, Leon2019]: dynamical symmetries, Cartan symmetries, and dynamical similarities. In this paper, we will focus on the characterization of such infinitesimal symmetries of contact Hamiltonian systems, providing new ways to compare and relate the aforementioned ones, and in some cases to reconstruct invariants of the dynamics.
Usually, vector fields on a contact manifold are decomposed into their projection to the contact distribution, known as the horizontal component, and the rest, known as vertical component. A different decomposition of the vector fields introduced in [projective, vectorcontactstructure, Zadra], instead, splits the vector fields into a horizontal component and an Hamiltonian one. This will be the central ingredient in comparing and characterizing the different symmetries.
Since this new construction is global and tied to the choice of contact form, we introduce an alternative description in terms of tensor densities which ends up being independent of it and providing an additional tool for the characterization.
We believe that the characterizations provided here could be useful tools for the study of the dynamical consequences of symmetries in contact Hamiltonian systems, especially once applied to study explicit examples of application and their contact integrability properties.
The paper is structured as follows. In Section 2 we briefly review contact Hamiltonian systems and introduce the necessary notation. In Section 3 we introduce and describe the horizontal–Hamiltonian decomposition and its expression in terms of tensor densities. In Section 4 we derive the characterization of symmetries and dissipated quantities of contact Hamiltonian systems. All the notions of symmetry are discussed in their separate subsections: Dynamical symmetries in Section 4.1, Dynamical similarities in Section 4.3, with a particular focus on the special case of Scaling symmetries in Section 4.2, and Cartan symmetries in Section 4.4. Finally, in Section 5, we begin to explore how this new characterization of symmetries can help to reconstruct constants of motion for contact Hamiltonian systems.
2 Contact Geometry
In this section, we will introduce the basic concepts of contact geometry, focusing on the aspects that are relevant for our analysis of contact Hamiltonian systems.
Definition 1.
A contact manifold $(M,\Delta)$ is a pair of a differentiable manifold $M$ of odd dimension $12n+$ 1 and a maximally non-integrable smooth distribution $\Delta\subset TM$ of co-dimension $1$. We say that the contact manifold is exact (or co-oriented [geiges]) if the distribution $\Delta$ coincides with the kernel of a $1$-form $\eta$ such that
$\eta\wedge(d\eta)^{\wedge n}\neq 0,$ (1)
where
$d\eta^{\wedge n}=\underbrace{d\eta\wedge\ldots\wedge d\eta}_{n-\text{times}}.$ (2)
We will call the distribution $\Delta$ the horizontal distribution, and its sections horizontal vector fields.
Remark 1.
With an abuse of notation, we will refer to $\Delta$ also as the space of the sections $\Gamma(\Delta)$ over the manifold $M$.
The distinguishing property of exact contact manifold is that we can globally describe the contact distribution by just considering one particular $1$-form, what is also referred to as the co-orientation of the contact distribution. This is not generally true: a local contact form always exists but it may not possible to extend it globally. In any case, including exact contact manifolds, the choice of the contact form is not unique: any other $1$-form that is proportional to the contact form by a non-vanishing real function is again a contact form that describes the same contact distribution, i.e. for all $f:M\to\mathbb{R}^{+}$
$\ker\eta=\ker f\eta.$ (3)
Once a contact form has been fixed, the $(2n+1)$-form in equation (1) is the natural choice of volume form for the contact manifold. Moreover, we can use the non-integrability condition (1) to describe the part of the tangent space that is transverse to the contact distribution. This transversal component, leads to the choice of a particular generator for the complement of the contact distribution: the Reeb vector field.
Proposition 1 ([geiges]).
Let $(M,\eta)$ an exact contact manifold. There exists a unique vector field $\mathcal{R}\equiv\mathcal{R}_{\eta}\in\Gamma(TM)$ such that
$\eta(\mathcal{R})=1,\quad\mbox{and}\quad d\eta(\mathcal{R},\cdot)=0.$ (4)
Such a vector field is called Reeb vector field for $\eta$.
It should be clear from the previous statement that the Reeb vector field $\mathcal{R}_{\eta}$ is tightly connected to the choice contact form $\eta$, and in general will not be as rigid as $\ker\eta$. Indeed, despite the simplicity of equation (3), the corresponding transformation law $\mathcal{R}_{\eta}\mapsto\mathcal{R}_{f\eta}$ for the Reeb vector field is not analogously nice and the flows of the two Reeb vector fields can be quite different.
To study in depth the transformation of contact forms and related quantities, we should introduce the notion of contactomorphism, that is, a transformation that preserves the contact distribution.
Definition 2.
Let $(M,\Delta_{M})$ and $(N,\Delta_{N})$ be two contact manifolds. A smooth map $\phi:M\to N$ is called a contactomorphism (or contact transformation) if
$\Delta_{M}\subseteq\phi^{*}(\Delta_{N}).$ (5)
This definition can be extended to exact contact manifolds by rewriting the condition (5) in terms of the contact form.
Proposition 2.
Let $(M,\eta_{M})$ and $(N,\eta_{N})$ be two exact contact manifolds. If $\phi:M\to N$ is a smooth contactomorphism, then
$\phi^{*}\eta_{N}=F_{\phi}\eta_{M},$ (6)
where $F_{\phi}\in C^{\infty}(M)$ is a non-vanishing smooth real function. Moreover, if $\phi^{t}:\mathbb{R}\times M\to N$, is a $1$-parameter group of smooth contactomorphisms, its infinitesimal generator $Y_{\phi}$ satisfies
$L_{Y_{\phi}}\eta=-f_{Y_{\phi}}\eta,$ (7)
where $f_{Y_{\phi}}=\frac{dF_{\phi^{t}}}{dt}\in C^{\infty}(M)$ is a smooth real function. We refer to $Y_{\phi}$ as infinitesimal contactomorphism.
Remark 2.
Let $(M,\eta)$ be an exact contact manifold. If a contactormorphism $g:M\to M$ acts on $\eta$ as
$g^{*}\eta=f_{g}\eta,$ (8)
by a non-vanishing function $f_{g}$, the volume form $\eta\wedge(d\eta)^{\wedge n}$ is transformed as
$g^{*}(\eta\wedge(d\eta)^{\wedge n})=(f_{g})^{n+1}\eta\wedge(d\eta)^{\wedge n}.$ (9)
2.1 Contact Hamiltonian systems
Any 1-parameter group of contactomorphisms is the flow of a contact Hamiltonian function [geiges], so we have already implicitly introduced contact Hamiltonian systems in Proposition 2. A more practical definition can be obtained by mimicking the construction of Hamiltonian vector fields in the symplectic setting [Bravetti2017, Leon2019].
Definition 3.
A contact Hamiltonian system is a triple $(M,\eta,\mathcal{H})$ where the pair $(M,\eta)$ is an exact contact manifold, and $\mathcal{H}\in C^{\infty}(M)$ is a real function called contact Hamiltonian. Any contact Hamiltonian induces a vector field, called the contact Hamiltonian vector field associated to $\mathcal{H}$, defined by the equations
$\eta(X_{\mathcal{H}})=-\mathcal{H},\qquad d\eta(X_{\mathcal{H}},\cdot)=d\mathcal{H}-\mathcal{R}(\mathcal{H})\eta.$ (10)
One of the main differences between the symplectic and the contact settings is that in the latter the Hamiltonian flow does not preserve the Hamiltonian function. Indeed, it follows immediately from (10) that
$X_{\mathcal{H}}(\mathcal{H})=-\mathcal{R}(\mathcal{H})\mathcal{H}.$ (11)
For a symplectic Hamiltonian system, the right hand side of (11) vanishes and, therefore, the energy (i.e., in this case, the Hamiltonian itself) is conserved along the Hamiltonian flow, see (161). Therefore, in the contact case, unless $\mathcal{H}=0$ or $\mathcal{R}(\mathcal{H})=0$, one can think of (11) as encoding a “non-conservation” of energy.
Remark 3.
Equation (10) and Cartan’s magic formula, imply that
$L_{X_{\mathcal{H}}}\eta=d\iota_{X_{\mathcal{H}}}\eta+\iota_{X_{\mathcal{H}}}d\eta=-\mathcal{R}(\mathcal{H})\eta\,.$ (12)
This is a very convenient identity when working with contact Hamiltonian vector fields, and is sometimes used as an alternative definition, in place of the right hand side of (10).
The two propositions that follow are important tools for studying infinitesimal symmetries of contact Hamiltonian systems. As pointed out already in Proposition 2, the action of an infinitesimal contactormorphism leaves the distribution invariant, and contact Hamiltonian vector fields are infinitesimal contactomorphisms. In fact, the converse is also true [geiges, Leon2019]: any infinitesimal contactomorphism is a contact Hamiltonian vector field in the sense of (10).
Proposition 3.
Contact Hamiltonian vector fields preserve the contact distribution, i.e. if $X_{f}$ is one such vector field, $[X_{f},\Delta]\subseteq\Delta$.
Proof.
Consider a contact Hamiltonian vector field $X_{g}$ and let $Y\in\Delta$ be an arbitrary vector field tangent to the contact distribution. Computing the Lie bracket, and observing that $\eta(Y)=0$, we get
$\eta(\left[X_{g},Y\right])$ $=d\eta(X_{g},Y)+Y(\eta(X_{g}))-X_{g}(\eta(Y))$ (13)
$=dg(Y)-R(g)\eta(Y)-dg(Y)$ (14)
$=0,$ (15)
therefore $[X_{g},Y]\in\Delta$. ∎
Moreover, the space of contact Hamiltonian vector fields is closed under the action of Lie brackets, as shown in the next proposition.
Proposition 4.
If $X_{g}$ and $X_{f}$ are contact Hamiltonian vector fields, then their Lie bracket $[X_{g},X_{f}]=X_{-\eta([X_{g},X_{f}])}$ is itself a contact Hamiltonian vector field.
Proof.
Let $h:=-\eta([X_{g},X_{f}])$. We can compute the Lie derivative $L_{\left[X_{g},X_{f}\right]}\eta$ in two ways by using the definition of a contact Hamiltonian vector field and equation (12):
$L_{\left[X_{g},X_{f}\right]}\eta=-dh+d\eta(\left[X_{g},X_{f}\right],\cdot),\qquad\mbox{and}$ (16)
$L_{\left[X_{g},X_{f}\right]}\eta=L_{X_{g}}L_{X_{f}}\eta-L_{X_{f}}L_{X_{g}}\eta=-\left(X_{g}(\mathcal{R}(f))-X_{f}(\mathcal{R}(g))\right)\eta.$ (17)
Thus, in particular, we have
$-\left(X_{g}(\mathcal{R}(f))-X_{f}(\mathcal{R}(g))\right)\eta+dh-d\eta(\left[X_{g},X_{f}\right],\cdot)=0.$ (18)
Contracting (18) with the Reeb vector field, we obtain
$-X_{g}(\mathcal{R}(f))+X_{f}(\mathcal{R}(g))+\mathcal{R}(h)=0.$ (19)
This can be used in (18) to get
$d\eta(\left[X_{g},X_{f}\right],\cdot)=dh-\mathcal{R}(h)\eta.$ (20)
In summary, if $Y:=[X_{g},X_{f}]$, we have shown that
$\eta(Y)=-h,\qquad\mbox{and}\qquad d\eta(Y,\cdot)=dh-\mathcal{R}(h)\eta,$ (21)
that is, $[X_{g},X_{f}]=X_{h}$ is the contact Hamiltonian vector field associated to $h$. ∎
In what follows, we also need to introduce a particular contact Hamiltonian system, whose contact Hamiltonian is the constant function $\mathcal{H}\equiv-1$.
Definition 4.
The Reeb system on a contact manifold $(M,\eta)$ is the contact Hamiltonian system defined by the triple $(M,\eta,\mathcal{H}\equiv-1)$. In the rest of the manuscript we will abuse notation and denote it simply $(M,\eta,-1)$.
At least locally, every contact Hamiltonian system can be straightened into a Reeb system [Bravetti2015]. Indeed, if the Hamiltonian of a contact Hamiltonian system $(M,\eta,\mathcal{H})$ is not vanishing at a point $x_{0}\in M$, then there is an open neighbourhood $V\subset M$ of $x_{0}$ on which $\mathcal{H}\neq 0$. In this neighbourhood we can define a contact transformation that reduces the contact Hamiltonian system to a Reeb system $(V,\tilde{\eta},-1)$: indeed, the $1$-form
$\tilde{\eta}\rvert_{V}:=-\frac{1}{\mathcal{H}}\eta\rvert_{V},$ (22)
defines on $V$ the same contact structure $\Delta=\ker\tilde{\eta}=\ker\eta$ as $\eta$. Moreover, the Hamiltonian vector field $X_{\mathcal{H}}$ is the Reeb vector field for $\tilde{\eta}$: this directly follows from the explicit computation
$\tilde{\eta}(X_{\mathcal{H}})=-\frac{\eta(X_{\mathcal{H}})}{\mathcal{H}}=1,\qquad d\tilde{\eta}(X_{\mathcal{H}},\cdot)=\left(+\frac{d\mathcal{H}\wedge\eta}{\mathcal{H}^{2}}-\frac{d\eta}{\mathcal{H}}\right)(X_{\mathcal{H}},\cdot)=0,$ (23)
and by uniqueness of the Reeb vector field, see Proposition 1.
Remark 4.
On the converse, a contact transformation $\psi:(M,\eta)\to(N,\eta^{\prime})$ whose pullback scales the contact form by a non-vanishing function $F_{\psi}:M\to\mathbb{R}_{+}$ as in Proposition 2, maps the Reeb vector field into a Hamiltonian vector field [geiges, Zadra]:
$\psi_{*}\mathcal{R}=-X^{\prime}_{F_{\psi}\circ\psi^{-1}},$ (24)
where for $X^{\prime}$ we refer to the contact Hamiltonian vector field computed from $\eta^{\prime}$.
2.2 Jacobi Structure on contact manifolds
The function $h=-\eta([X_{g},X_{f}])$ introduced in the proof of Proposition 4 has a deeper meaning in the context of contact Hamiltonian systems. It represents the natural Jacobi bracket structure on the functions on a contact manifold [Leon2019, de_Le_n_2020].
Definition 5 ([Kirillov1976]).
A Jacobi manifold is a smooth manifold $M$ endowed with a Jacobi structure, that is, a pair of a skew bivector field (i.e., a skew-symmetric, contravariant 2-tensor on $TM$) $\Lambda(\cdot,\cdot)$ and a vector field $\mathcal{E}\in\Gamma(TM)$ such that
$[\Lambda,\mathcal{E}]_{SN}$ $=0,$ (25)
$[\Lambda,\Lambda]_{SN}$ $=-2\Lambda\wedge\mathcal{E},$ (26)
where $[\cdot,\cdot]_{SN}$ denotes the Schouten-Nijenhuis brackets [Marsden2010].
A Jacobi structure on a manifold induces a bilinear map between functions that are continuously differentiable; we call this map Jacobi bracket.
Definition 6.
If $(M,\Lambda,\mathcal{E})$ is a Jacobi manifold, there is a natural product structure on smooth functions called Jacobi bracket and defined by
$\{f,g\}_{J}=\Lambda(df,dg)-g\mathcal{E}(f)+f\mathcal{E}(g).$ (27)
Remark 5.
If $\mathcal{E}=0$, then the structure induced by $\Lambda(\cdot,\cdot)$ is a Poisson structure [Kirillov1976].
Remark 6.
The bivector $\Lambda$ induces a morphism of vector bundles $T^{*}M\to T^{**}M\simeq TM$ via the mapping $\nu\mapsto\Lambda(\nu,\cdot)$. This will be very handy in the constructions that follow.
Let us revisit this in the contact setting. The choice of a contact form on an exact contact manifold, induces a Jacobi structure on the manifold and thus a Jacobi bracket on smooth functions.
The non-integrability of the contact structure on an exact contact manifold $(M,\eta)$ allows one to define an isomorphism of vector bundles
$\flat$ $:TM\to T^{*}M$ (28)
$\quad X\mapsto\iota_{X}\,d\eta+\eta(X)\eta,$ (29)
and a corresponding isomorphism of $C^{\infty}(M)$-modules between vector fields and one-forms that we will denote by the same symbol [de_Le_n_2020, Section 2.2]. This then allows to construct a natural Jacobi structure on the manifold as follows.
Lemma 1 ([de_Le_n_2020]).
On a contact manifold $(M,\eta)$, there is a natural Jacobi structure defined by the contact form $\eta$ as follows:
$\Lambda(\sigma,\nu)=-d\eta(\flat^{-1}\sigma,\flat^{-1}\nu),\quad\mbox{and}\quad\mathcal{E}=-\mathcal{R}\mbox{ is the Reeb vector field}.$ (30)
Remark 7.
On an exact contact manifold $(M,\eta)$, we denote by {,}$\{\cdot,\cdot\}_{\eta}:=\{\cdot,\cdot\}_{J}$ the Jacobi brackets induced by the contact form.
Similarly to the symplectic case, the contact Hamiltonian vector field has an equivalent expression in terms of the Jacobi structure.
Lemma 2.
Let $(M,\eta)$ be a contact manifold and $(\Lambda,\mathcal{E})$ be the corresponding Jacobi structure. Then, the contact Hamiltonian vector field $X_{f}$ associated to $f$ has the form
$X_{f}=\Lambda(df,\cdot)+f\mathcal{E}=\Lambda(df,\cdot)-f\mathcal{R}.$ (31)
Moreover, for any $f,g\in C^{\infty}(M)$, the Jacobi bracket $\{f,g\}_{\eta}$ can be rewritten in terms of the contact structure as
$\{f,g\}_{\eta}=X_{f}(g)+g\mathcal{R}(f).$ (32)
Proposition 4, then, implies that we can relate the Lie bracket of two contact Hamiltonian vector fields to the Jacobi brackets of their Hamiltonian functions: between the set of the contact Hamiltonian vector fields and the contact Hamiltonian functions there is an anti-isomorphism of Lie algebras.
Proposition 5.
Let $X_{f}$ and $X_{g}$ two Hamiltonian vector fields. Then,
$\eta([X_{f},X_{g}])=-\{f,g\}_{\eta},$ (33)
and
$\left[X_{f},X_{g}\right]=X_{\left\{f,g\right\}_{\eta}}.$ (34)
That is, the Lie bracket of these two vector fields is a contact Hamiltonian vector field with Hamiltonian function $\{f,g\}_{\eta}$.
Proof.
The first equation is proven in [de_Le_n_2020, Lemma 3]. Then Proposition 4 implies
$\left[X_{f},X_{g}\right]=X_{-\eta([X_{f},X_{g}])}=X_{\left\{f,g\right\}_{\eta}}.$ (35)
3 Decompositions of vector fields on contact manifolds
On an exact contact manifold $(M,\eta)$, the contact form $\eta$ naturally induces a splitting of the tangent space $T_{x}M$ at any point $x\in M$ into a component tangential to the horizontal distribution and the rest [geiges]. This local horizontal–vertical decomposition can be naturally extended to vector fields over $M$ with the aid of the Reeb vector field, since in this case $TM=\mathrm{span}\{\mathcal{R},\Delta\}$ holds globally.
Proposition 6.
Let $(M,\eta)$ a contact manifold, the $C^{\infty}(M)$ module of vector fields on $M$ admits the decomposition
$\Gamma(TM)=\ker\eta\oplus\ker d\eta.$ (36)
In view of equation (36), given $X\in\Gamma(TM)$, we can decompose it into
$X=h_{X}+\eta(X)\mathcal{R};$ (37)
where $h_{X}\in\Delta=\ker\eta$ and $\eta(X)$ is a smooth function. The vector field $\eta(X)\mathcal{R}$ is called the vertical part of $X$.
For our purposes, it will be convenient to consider a different decomposition, introduced in [projective, vectorcontactstructure, Zadra], that replaces the vertical component with a contactomorphism.
3.1 The Hamiltonian–horizontal decomposition
The Hamiltonian–horizontal decomposition splits vector fields on a contact manifolds in a component described by a contact Hamiltonian vector field and a horizontal one.
Proposition 7.
Let $(M,\eta)$ be a contact manifold. Any vector field $\xi\in\Gamma(TM)$ admits a unique decomposition
$\xi=X_{\phi_{\xi}}+\delta_{\xi},$ (38)
where $\phi_{\xi}=\ -\eta(\xi)\in C^{\infty}(M)$ and $\delta_{\xi}\in\Delta=\ker\eta$.
Proof.
Equation (10) applied to $\delta_{\xi}=\xi-X_{\phi_{\xi}}$ implies
$\eta(\xi-X_{\phi_{\xi}})=\eta(\xi)-\eta(X_{-\eta(\xi)})=0.$ (39)
That is, $\delta_{\xi}$ a horizontal vector field.
Since $\eta(\xi)$ is a well-defined smooth function on $M$, its Hamiltonian vector field $X_{\phi_{\xi}}$ is uniquely defined, and therefore the horizontal vector field $\delta_{\xi}=\xi-X_{\phi_{\xi}}$ is also unique. ∎
Differently from the horizontal–vertical decomposition, which is inherently local, the Hamiltonian–horizontal decomposition is necessarily a global construction: the set of Hamiltonian vector fields evaluated at a point can span the entire tangent space. While this can be seen as a disadvantage, this construction is crucial to characterize the relations between the various symmetries of contact Hamiltonian systems.
Remark 8.
In the study of symmetries, the Hamiltonian–horizontal decomposition is preferred to the usual horizontal–vertical one (36) because it behaves better with the Lie brackets, as exemplified by Propositions 2 and 3.
In practice, the relation between a vector field $\xi$ and the contact Hamiltonian system $(M,\eta,\mathcal{H})$ is often characterised by their Lie bracket $[\xi,X_{\mathcal{H}}]$. The Hamiltonian–horizontal decomposition, reduces the computation to
$[\xi,X_{\mathcal{H}}]=[X_{\phi_{\xi}},X_{\mathcal{H}}]+[\delta_{\xi},X_{\mathcal{H}}].$ (40)
Here, the first term $[X_{\phi_{\xi}},X_{\mathcal{H}}]$ is a Hamiltonian vector field, by Propositions 4 and 5, and the second is horizontal, since by (12)
$\eta([\delta_{\xi},X_{\mathcal{H}}])=\iota_{\delta_{\xi}}L_{X_{\mathcal{H}}}\eta-L_{X_{\mathcal{H}}}\iota_{\delta_{\xi}}\eta=-\mathcal{R}(\mathcal{H})\iota_{\delta_{\xi}}\eta-L_{X_{\mathcal{H}}}\iota_{\delta_{\xi}}\eta=0.$ (41)
This means that the sum in equation (40) is manifestly the Hamiltonian decomposition of the commutator, i.e.,
$[\xi,X_{\mathcal{H}}]=X_{-\{\phi_{\xi},\mathcal{H}\}_{\eta}}+\delta_{[\xi,X_{\mathcal{H}}]}.$ (42)
One may still wonder what is the relation between the Hamiltonian–horizontal decomposition and the horizontal–vertical one from (36).
Remark 9.
Since the Reeb vector field is orthogonal to the distribution, any vertical vector field $Y$ on a contact manifold $(M,\eta)$ has to be parallel to it, and thus by (31), it can be written as
$Y=f\mathcal{R}=X_{-f}+\Lambda(df,\cdot),$ (43)
where $\Lambda(\cdot,\cdot)$ is the skew bivector field of the Jacobi structure induced by $\eta$ as in Lemma 1.
Therefore, a vertical vector field has Hamiltonian–horizontal decomposition $Y=X_{\phi_{Y}}+\delta_{Y}$ with
$X_{\phi_{Y}}=-X_{f},\quad\mbox{and}\quad\delta_{Y}=+\Lambda(df,\cdot).$ (44)
Here we want to stress the key role of the Jacobi structure in extracting the Hamiltonian component from the vertical field.
This centrality of the Jacobi structure will keep coming back also in the upcoming sections. By providing a natural way to identify functions in involution, symmetries, and dissipated quantities, it will play an important role in the study of the symmetries of contact Hamiltonian systems.
3.2 Hamiltonian–horizontal decomposition and tensor densities
The Hamiltonian–horizontal decomposition presented in the previous section strictly depends on the choice of the contact form. If we consider a contactomorphism $\eta\mapsto f\eta$ as in Proposition 2, the Hamiltonian function will be rescaled by $f$ and a more complicated transformation law will hold for $\delta_{Y}$. It is possible to relax this dependence on the contact form by introducing tensor densities, as done in [projective, Chapter 7.5] and [vectorcontactstructure].
We postpone the technical definitions to Appendix A and just recall here the relevant consequences in the case of exact contact manifolds. In particular, what matters is equation (45): on $(M,\eta)$ a tensor density $\phi$ of degree $\lambda$ is an object with the local expression
$\phi=f(x)\ \eta^{(n+1)\lambda}.$ (45)
In particular, this means that the choice of a contact form $\eta$ allows to express tensor densities (as a module over the group of contactormorphisms) not in terms of the the volume form induced by $\eta$, but directly in terms of $\eta$. In particular, it turns out that contactomorphisms are a very special case of tensor densities.
Proposition 8 ([projective, vectorcontactstructure]).
The set of infinitesimal contactormorphisms is isomorphic to the set of tensor densities of degree $-\frac{1}{n+1}$.
This means that we can use equation (45) to identify the Hamiltonian of a contact Hamiltonian system $(M,\eta,\mathcal{H})$ with a tensor density $\psi_{X_{\mathcal{H}}}$ of degree $-\frac{1}{n+1}$, which is given explicitly by
$\psi_{X_{\mathcal{H}}}=\eta(X_{\mathcal{H}})\ \eta^{-1}=-\mathcal{H}\ \eta^{-1},$ (46)
where to make sense of $\eta^{-1}$ we view it as a tensor density of degree $-\frac{1}{n+1}$. This representation is useful for two reasons [projective, vectorcontactstructure]:
  • first, the expression of the contact Hamiltonian on the right hand side of (46) is independent of the choice of the contact form;
  • second, the Jacobi brackets of two contact Hamiltonian functions are the Poisson brackets of the corresponding densities.
This directly motivates the following definition.
Definition 7.
A tensor density $\sigma$ is an invariant tensor density, if it is invariant under contactomorphism [vectorcontactstructure].
To find an invariant representation of horizontal vector fields of $(M,\eta)$ in terms of densities, it is convenient to introduce the set $\Omega_{h}^{2}(M)$ of 2-forms vanishing on the contact distribution, that is, 2-forms that can be expressed as $\eta\wedge\beta$ with $\beta\in\Omega^{1}(M)$.
Proposition 9 ([projective, vectorcontactstructure]).
Let $(M,\eta)$ be an exact contact manifold. The set $\Delta=\ker\eta$ of horizontal vector fields is isomorphic to
$\Omega_{h}^{2}(M)\otimes\mathcal{F}_{-\frac{2}{n+1}}(M),$ (47)
where $\mathcal{F}_{-\frac{2}{n+1}}$ denotes the set of tensor densities on $M$ of degree $-\frac{2}{n+1}$.
With Propositions 8 and 9 at hand, we can assign to each vector field two contact-invariant densities which represents, respectively, its horizontal and Hamiltonian parts.
Proposition 10.
Let $(M,\eta)$ be an exact contact manifold. The Hamiltonian–horizontal decomposition $\xi=X_{\phi_{\xi}}+\delta_{\xi}$ of any vector field $\xi\in\Gamma(TM)$ has an expression in terms of invariant tensor densities of the form $\psi_{\xi}+\sigma_{\xi}$, where
  • the Hamiltonian component $\psi_{\xi}=\eta(\xi)\ \eta^{-1}$ is a tensor density of degree $-\frac{1}{n+1}$;
  • and, the horizontal component $\sigma_{\xi}$ is the tensor density of degree $-\frac{2}{n+1}$ given by
    $\sigma_{\xi}=-\eta\wedge d\eta\left(Y+X_{\eta}(Y),\cdot\right)\otimes\eta^{-\frac{2}{n+1}}.$ (48)
4 Symmetries of contact Hamiltonian systems and their decomposition
In the symplectic setting, any function that is in involution with the Hamiltonian with respect to the Poisson brackets, is a constant of motion of the system and the Hamiltonian vector fields generated by functions that are in involution are pairwise commuting [de_Le_n_2020, Leon2019].
For contact Hamiltonian systems, Proposition 5 tells us that the latter property remains true. However, the quantities that give rise to symmetries are not necessarily conserved along the contact Hamiltonian flow: we already saw with equation (11) that the contact Hamiltonian itself is in general not preserved along its own contact flow. This is why, symmetries of contact Hamiltonian systems have been introduced in terms of “dissipated” quantities [de_Le_n_2020, Leon2019].
Definition 8.
Let $(M,\eta,\mathcal{H})$ be a contact Hamiltonian system. A function $f\in C^{\infty}(M)$ is a dissipated quantity if
$\{f,\mathcal{H}\}_{\eta}=0.$ (49)
That is, if $f$ is in involution with the contact Hamiltonian $\mathcal{H}$ with respect to the Jacobi brackets induced by the contact form $\eta$.
A conserved quantity can nevertheless be recovered by considering ratios of dissipated quantities: if we consider equation (49) and (32), we can observe that any two quantities that are in involution, will dissipate with the same ratio.
Proposition 11 ([Leon2019]).
Let $(M,\eta,\mathcal{H})$ be a contact Hamiltonian system. If $f,g$ are two function in involution with the Hamiltonian $\mathcal{H}$, their ratio
$\kappa=\frac{f}{g}$ (50)
is conserved along the flow of $\mathcal{H}$.
In what follows, we will study other types of infinitesimal symmetries that are not necessarily Hamiltonian, and thus, not necessarily infinitesimal contact transformation: dynamical symmetries, Cartan symmetries (both introduced on the contact setting in [de_Le_n_2020]) and dynamical similarities [Bravetti2023, Sloan2018, Sloan2021].
Definition 9.
Let $(M,\eta,\mathcal{H})$ a contact Hamiltonian system. A vector field $Y\in\Gamma(TM)$ is called
  • dynamical symmetry if
    $\eta\left(\left[Y,X_{\mathcal{H}}\right]\right)=0\,;$ (51)
  • dynamical similarity if there exists $\lambda\in C^{\infty}(M)$, such that
    $\left[Y,X_{\mathcal{H}}\right]=\lambda X_{\mathcal{H}}\,.$ (52)
  • Cartan symmetry if there exist $a,g\in C^{\infty}(M)$ such that
    $\begin{cases}L_{Y}\eta=a\eta+dg\\ Y(\mathcal{H})=a\mathcal{H}+gR(\mathcal{H})\end{cases}\,.$ (53)
The following theorem shows that, at least for the first two kinds of symmetries above, it is possible to recover “dissipated” quantities. This is particularly interesting since the symmetry $Y$ is, in general, not a contact Hamiltonian vector field.
Theorem 1 ([de_Le_n_2020]).
Let $(M,\eta,\mathcal{H})$ a contact Hamiltonian system. Let $Y\in\Gamma(TM)$, then the following hold.
  • $Y$ is a dynamical symmetry if and only if $\{\eta(Y),\mathcal{H}\}_{\eta}=0$.
  • $Y$ is a Cartan symmetry if and only if $\{\eta(Y)-g,\mathcal{H}\}_{\eta}=0$, where $g$ is defined in (53).
Moreover $Y$ is both a dynamical symmetry and a Cartan symmetry if and only if $dg=0$ and $Y$ is a contact Hamiltonian vector field.
This statement already shows that there are nontrivial relations between these various different notions of symmetry. In the following subsections we will try to characterize them more finely in terms of their decompositions.
4.1 Dynamical symmetries
Dynamical symmetries are vector fields that commute with the contact Hamiltonian vector field modulo $\eta$, in the sense that their commutator with the contact Hamiltonian vector field belongs to the contact distribution. The Hamiltonian–horizontal decomposition is particularly well suited to capture this.
Proposition 12.
Let $(M,\eta,\mathcal{H})$ be a contact Hamiltonian system, and $Y\in\Gamma(TM)$. Then $Y$ is a dynamical symmetry for the system if and only if its Hamiltonian–horizontal decomposition $Y=X_{\phi_{Y}}+\delta_{Y}$ satisfies
$\{\phi_{Y},\mathcal{H}\}_{\eta}=0\,.$ (54)
Proof.
If $Y$ is a dynamical symmetry, this follows immediately from the first statement of Theorem 1 after observing that $\phi_{Y}=-\eta(Y)$, see Proposition 7.
Conversely, let $Y=X_{\phi_{Y}}+\delta_{Y}$. Then
$\eta([Y,X_{\mathcal{H}}])=\eta([X_{\phi_{Y}}+\delta_{Y},X_{\mathcal{H}}])=\eta([X_{\phi_{Y}},X_{\mathcal{H}}])=\{\eta(Y),\mathcal{H}\}_{\eta}=0.$ (55)
Which by the first statement of Theorem 1 concludes the proof. ∎
This means that $\phi_{Y}=-\eta(Y)$ is a dissipated quantity for the contact Hamiltonian system, and therefore by Proposition 10 we get the following corollary.
Corollary 1.
The invariant tensor density associated to a dynamical similarity is
$\mathfrak{K}=-\phi_{Y}\ \eta^{-1}.$ (56)
Proposition 12 makes it clear that the definition of dynamical symmetries restricts only the Hamiltonian part of the vector field, leaving the horizontal part free: roughly speaking, this reflects the fact that equation (51) provides only one condition out of the $12n+$ 1 required to specify a vector field on a contact manifold.
4.2 Scaling symmetries
The next symmetries that we want to investigate are dynamical similarities. We first focus on a specific family, the so-called scaling symmetries. A vector field $Y\in\Gamma(TM)$ is a scaling symmetry if it satisfies the equation (52) with $\lambda$ constant. Furthermore, the distribution spanned by $X_{\mathcal{H}}$ and $Y$ is involutive, so integrable by Frobenius theorem.
Proposition 13.
If $Y\in\Gamma(TM)$ is a scaling symmetry of the contact Hamiltonian system $(M,\eta,\mathcal{H})$, then $Y$ is of the form
$Y=X_{\phi_{Y}}+\delta_{Y},$ (57)
with $\phi_{Y}=-\eta(Y)$ and $[\delta_{Y},X_{\mathcal{H}}]=0$.
Proof.
By Definition 9, we have
$\left[Y,X_{\mathcal{H}}\right]=\lambda X_{\mathcal{H}}=X_{\lambda\mathcal{H}}.$ (58)
Since the right-hand side of equation (58) is Hamiltonian the left-hand side also has to be Hamiltonian. From Propositions 3, 4, and 7, we find that $\left[Y,X_{\mathcal{H}}\right]=\left[X_{\phi_{Y}},X_{\mathcal{H}}\right]+\left[\delta_{Y},X_{\mathcal{H}}\right]$, in particular, this means that for Proposition 3
$0=\delta_{\left[Y,X_{\mathcal{H}}\right]}=\left[\delta_{Y},X_{\mathcal{H}}\right].$ (59)
Corollary 2.
A scaling symmetry $Y\in\Gamma(TM)$ of a contact Hamiltonian system $(M,\eta,\mathcal{H})$, is purely Hamiltonian, i.e., if $\delta_{Y}=0$, if and only if $L_{Y}\eta=-\mathcal{R}(\phi_{Y})\eta$. In any case, the Hamiltonian vector field generated by the function $\phi_{Y}$ is a scaling symmetry of the system.
As a consequence of Proposition 13 and Remark 2, we can explicitly compute the action of the scaling symmetry $\tilde{Y}=X_{-\eta(Y)}$ on the Hamiltonian and the contact from:
$\begin{cases}\tilde{Y}(\mathcal{H})=\mathcal{H}\left(\lambda-\mathcal{R}(\phi_{Y})\right)\\ L_{\tilde{Y}}\eta=-\mathcal{R}(\phi_{Y})\eta\end{cases}\,.$ (60)
In [Bravetti2023a] the definition of scaling symmetry $W$ on a symplectic Hamiltonian system $(N,\omega,\mathfrak{H})$, with $(N,\omega)$ a symplectic manifold (cfr. Appendix B), is given by
$\begin{cases}W(\mathfrak{H})=\Lambda\mathfrak{H}\\ L_{W}\omega=\omega\end{cases}.$ (61)
Then, $W$ is a scaling symmetry for $(N,\omega,\mathfrak{H})$ if and only if
$\left[W,X_{\mathfrak{H}}\right]=\left(\Lambda-1\right)X_{\mathfrak{H}},$ (62)
where $X_{\mathfrak{H}}$ represent the symplectic Hamiltonian vector field. The proof follows immediately from the computation of the Lie derivative of $\omega$ with respect to $[W,X_{\mathfrak{H}}]$.
The main differences between (61) and (60) are: the appearance, in the latter, of the Reeb vector field acting on the Hamiltonian, and the fact that the scaling symmetry in the symplectic setting is also a Liouville vector field [Bravetti2023a, sympgeo] (see Definition 11).
Remark 10.
Equation (60) also allows one to express the adjoint representation of the algebra generated by commutators of $\phi_{Y}$ and $\mathcal{H}$ in terms of the square matrices
$\mathcal{H}=\begin{pmatrix}0&0\\ -1&0\end{pmatrix}\,,\qquad\phi_{Y}=\begin{pmatrix}0&0\\ 0&\lambda\end{pmatrix}\,,$ (63)
whose commutator reproduces the action of the Lie brackets.
One can take advantage of such matrix representation and Definition 9, to find an explicit expression for the general element $\mathcal{X}$ of this two-dimensional subalgebra by means of the Baker-Campbell-Hausdorff formula [Hall2015]:
$\mathcal{X}=-\frac{t\tau\lambda}{1-\exp{(\tau\lambda)}}\mathcal{H}+\tau\ \phi_{Y},\qquad\tau,t\in\mathbb{R}.$ (64)
These equations are even more relevant in the case our hamiltonian function remains bounded away from zero. There, we can rescale the Hamiltonian to a Reeb system, as we have already seen in Section 2.1, and the scaling symmetries simply further.
Corollary 3.
For a Reeb system $(M,\eta,-1)$, the scaling symmetries take the form
$\mathcal{R}(\phi_{Y})=-\left\{\phi_{Y},-1\right\}_{\eta}=-\eta([Y,X_{-1}])=\lambda,$ (65)
and therefore one can define the quantity
$\kappa=\mathcal{R}(\phi_{Y}),$ (66)
whose value remains constant along the flow of the Hamiltonian. In particular, this means that $-\kappa$ is a dissipated quantity.
We can use this result to find dissipated quantities for general contact Hamiltonian systems.
Proposition 14.
Let $(M,\eta,\mathcal{H})$ be a contact Hamiltonian system with a scaling symmetry $Y$. Then
$\Phi=-\frac{\phi_{Y}}{\mathcal{H}},$ (67)
is constant along the flow $X_{\mathcal{H}}(\Phi)=-\lambda\in\mathbb{R}$. Therefore, $\tilde{\Phi}\mathcal{H}$ is a dissipated quantity.
Proof.
This follows by explicit computation
$X_{\mathcal{H}}(\Phi)=X_{\mathcal{H}}\left(-\frac{\phi_{Y}}{\mathcal{H}}\right)=\frac{-X_{\mathcal{H}}(\phi_{Y})\mathcal{H}+\phi_{Y}X_{\mathcal{H}}(\mathcal{H})}{\mathcal{H}^{2}}=-\lambda.$ (68)
Example 1.
Consider the contact Hamiltonian system $\left(\mathbb{R}^{3},ds-pdq,\mathcal{H}(q,p,s)=p\right)$.
The function $\mathfrak{f}(q,p,s)=qp$ is the Hamiltonian function of a scaling symmetry with parameter $\lambda=1$. Indeed,
$\{p,qp\}_{\eta}=-p.$ (69)
By Proposition 14, we find that the function $\Phi(q,p,s)=q$ has a constant flow.
4.3 Dynamical similarities
When $\lambda$ is not a constant but a genuine function, the result of Proposition 13 no longer holds. Indeed, dynamical similarities in general have non-vanishing Hamiltonian and horizontal components. To prove this we need a small preliminary Lemma.
Lemma 3.
Let $f,g$ be two smooth functions on an exact contact manifold $(M,\eta)$, then
$X_{fg}=fX_{g}+gX_{f}+fg\,\mathcal{R}.$ (70)
Proof.
Contracting $\eta$ by each side of equation (70) we obtain
$\begin{cases}\eta(X_{fg})=-fg,\\ \eta(fX_{g}+gX_{f}+fg\,\mathcal{R})=f\eta(X_{g})+g\eta(X_{f})+fg=-fg.\end{cases}$ (71)
While the corresponding contraction of $d\eta$ gives
$\begin{cases}d\eta(X_{fg},\cdot)=d(fg)-\mathcal{R}(fg)\eta,\\ d\eta(fX_{g}+gX_{f}+fg\,\mathcal{R},\cdot)=f(dg-\mathcal{R}(g)\eta)+g(df-\mathcal{R}(f)\eta)=d(fg)-\mathcal{R}(fg)\eta.\end{cases}$ (72)
This shows that both sides of the equation define the same contact Hamiltonian vector field, thus concluding the proof. ∎
Recalling Remark 9, the following result should not come completely as a surprise.
Proposition 15.
Let $Y=X_{\phi_{Y}}+\delta_{Y}$ be the Hamiltonian–horizontal decomposition of a dynamical similarity of $(M,\eta,\mathcal{H})$. Then there exists $\lambda\in C^{\infty}(M)$ such that
$\{\mathcal{H},\phi_{Y}\}_{\eta}=-\lambda\mathcal{H}\qquad\text{and}\qquad[\delta_{Y},X_{\mathcal{H}}]=-\mathcal{H}\Lambda(d\lambda,\cdot).$ (73)
Proof.
From equation (52), we recover
$\eta\left(\left[Y,X_{\mathcal{H}}\right]\right)$ $=-\lambda\mathcal{H},$ (74)
which reads
$\eta([X_{\phi_{Y}},X_{\mathcal{H}}])+\eta([\delta_{Y},X_{\mathcal{H}}])$ $=-\lambda\mathcal{H}$ (75)
after expanding $Y=X_{\phi_{Y}}+\delta_{Y}$ in its Hamiltonian–horizontal decomposition. In the same way as (41), $\eta([\delta_{Y},X_{\mathcal{H}}])=0$, therefore
$-\lambda\mathcal{H}=\eta([X_{\phi_{Y}},X_{\mathcal{H}}])=\iota_{X_{\phi_{Y}}}L_{X_{\mathcal{H}}}\eta-L_{X_{\mathcal{H}}}\iota_{X_{\phi_{Y}}}\eta.$ (76)
Equation (12), the definition of contact Hamiltonian vector fields and equation (32) then imply
$-\lambda\mathcal{H}=\mathcal{R}(\mathcal{H})\phi_{Y}+X_{\mathcal{H}}(\phi_{Y})=\{\mathcal{H},\phi_{Y}\}_{\eta}.$ (77)
Therefore, starting from the definition of scaling symmetry, one can use Proposition 4 and Lemma 3 to show
$\lambda X_{\mathcal{H}}=[Y,X_{\mathcal{H}}]=X_{-\eta([Y,X_{H}])}+[\delta_{Y},X_{H}]=X_{\lambda\mathcal{H}}-\mathcal{H}X_{\lambda}-\lambda\mathcal{H}\mathcal{R}.$ (78)
Comparing terms in the last identity for the Hamiltonian–horizontal decomposition above, we get
$[\delta_{Y},X_{\mathcal{H}}]=-\mathcal{H}X_{\lambda}-\mathcal{H}\lambda\mathcal{R}=-\mathcal{H}\,\Lambda(d\lambda,\cdot),$ (79)
which concludes the proof. ∎
The local straightening of a contact Hamiltonian system to a Reeb system can simplify computations by providing an explicit function for $\lambda$.
Corollary 4.
If $Y\in\Gamma(TM)$ is a dynamical similarity of the Reeb system $(M,\eta,-1)$, then $\lambda=-\mathcal{R}(\phi_{Y})$.
Proof.
From proposition 15 follows that
$\{\phi_{Y},-1\}_{\eta}=\lambda,$ (80)
from which $\lambda=-\mathcal{R}(\phi_{Y})$. ∎
And, moreover, one can show that the dynamical similarity of a Reeb system is preserved by contactomorphisms.
Proposition 16.
Let $Y\in\Gamma(TM)$ be a dynamical similarity of $(M,\eta,-1)$. Let $\psi:M\to M$ be a contactomorphism such that $\psi^{*}\eta^{\prime}=f_{\psi}\eta$. Then $\psi_{*}Y$ is a dynamical similarity for $(M,\eta^{\prime},-f_{\psi}\circ\psi^{-1})$.
Proof.
Lie brackets commute with pushforwards, thus
$\psi_{*}[Y,\mathcal{R}]=[\psi_{*}Y,\psi_{*}\mathcal{R}]=\lambda\circ\psi^{-1}\psi_{*}\mathcal{R}.$ (81)
By Proposition 2 (and Remark 4), if $\psi^{*}\eta^{\prime}=f_{\psi}\eta$, then
$\psi_{*}\mathcal{R}=-X^{\prime}_{f_{\psi}\circ\psi^{-1}},$ (82)
where by $X^{\prime}_{f}$ we mean the Hamiltonian vector field induced on $(M,\eta^{\prime})$ by the function $f:M\to\mathbb{R}$. Then, if we denote $\tilde{Y}=\psi_{*}Y$, we get
$[\tilde{Y},X_{-f_{\psi}\circ\psi^{-1}}]=\lambda\circ\psi^{-1}X_{-f_{\psi}\circ\psi^{-1}}$ (83)
that is again a dynamical similarity of the system $(M,\eta^{\prime},-f_{\psi}\circ\psi^{-1})$. ∎
Of course, this result can be extended locally to any contact Hamiltonian system.
Proposition 17.
Let $Y\in\Gamma(TM)$ be a dynamical similarity of the contact Hamiltonian system $(M,\eta,\mathcal{H})$ such that $[Y,X_{\mathcal{H}}]=\lambda X_{\mathcal{H}}$ for some $\lambda\in C^{\infty}(M)$. If for any $x_{0}\notin\mathcal{H}^{-1}(0)$, there exist a neighborhood $U$ of $x_{0}$ and a contactomorphism
$\psi:(U\subset M,\eta)\to(\psi(U),\eta^{\prime}),$ (84)
such that $(\psi(U),\eta^{\prime}=-\frac{1}{\mathcal{H}}\eta,-1)$ is a Reeb system, then
$\lambda=X_{\mathcal{H}}\left(-\frac{\phi_{Y}}{\mathcal{H}}\right).$ (85)
Proof.
Let us consider the Reeb system $(\psi(U),\eta^{\prime}=-\frac{1}{\mathcal{H}}\eta,-1)$. By Proposition 16, it has a dynamical symilarity $\psi_{*}Y$ that satisfies $[\psi_{*}Y,\mathcal{R}^{\prime}]=\lambda_{\mathcal{R}^{\prime}}\mathcal{R}^{\prime}$, where $\mathcal{R}^{\prime}$ is the Reeb associated to $\eta^{\prime}$. By Corollary 4, $\lambda_{\mathcal{R}^{\prime}}=\mathcal{R}^{\prime}(\phi_{Y_{\mathcal{R}}})$. Applying the pushforward of the inverse $\psi^{-1}_{*}$, we map
$\lambda_{\mathcal{R}}\circ\psi=\lambda_{\mathcal{H}}.$ (86)
By combining the two
$d(\phi_{Y_{\mathcal{R}}}\circ\psi^{-1})(\psi^{-1}_{*}\mathcal{R})=X_{\mathcal{H}}\left(\phi_{Y_{\mathcal{R}}}\right).$ (87)
Finally by Proposition 2, we obtain
$\phi_{Y_{\mathcal{R}}}=-\frac{\phi_{Y_{\mathcal{H}}}}{\mathcal{H}},$ (88)
concluding the proof. ∎
4.4 Cartan symmetries
The definition of Cartan symmetries is more involved, and it is harder to give a direct geometric interpretation of the conditions in (53). Here the Hamiltonian–horizontal decomposition helps to clarify the role of the auxiliary function $g$.
Proposition 18.
Let $(M,\eta,\mathcal{H})$ be a contact Hamiltonian system, and $Y\in\Gamma(TM)$. Then $Y$ is a Cartan symmetry if and only if its Hamiltonian–horizontal decomposition has the form
$Y=X_{\phi_{Y}}+\Lambda(dg,\cdot),$ (89)
where $\Lambda$ is the skew-bivector field defining the natural Jacobi structure on $(M,\eta)$, and
$\{\phi_{Y}+g,\mathcal{H}\}_{\eta}=0,\quad\mbox{and}\quad a=-\mathcal{R}(\phi_{Y}+g).$ (90)
Proof.
Consider a Cartan symmetry $Y\in\Gamma(TM)$ of the contact Hamiltonian system $(M,\eta,\mathcal{H})$. It follows from Theorem 1 that $\eta(Y)+g$ is in involution with $\mathcal{H}$, that is, $\{-\phi_{Y}+g,\mathcal{H}\}_{\eta}=0$.
It remains to prove that, in this case, $Y$ decomposes as in equation (89). We know by Proposition 7 that, in general, $Y=X_{\phi_{Y}}+\delta_{Y}$. Then,
$a\eta+dg=L_{Y}\eta=L_{X_{\phi_{Y}}+\delta_{Y}}\eta=-\mathcal{R}(\phi_{Y})\eta+d\eta(\delta_{Y},\cdot),$ (91)
where for the left-hand side we used the definition of Cartan symmetry and for the right-hand side (12) and the fact that $\delta_{Y}$ is horizontal. Contracting by the Reeb vector field, we get that $a=-\mathcal{R}(\phi_{Y}+g)$.
Plugging the value of $a$ into equation (91) then implies
$d\eta(\delta_{Y},\cdot)=dg-\mathcal{R}(g)\eta=\iota_{X_{g}}d\eta=\iota_{\Lambda(dg,\cdot)-g\mathcal{R}}d\eta=d\eta(\Lambda(dg,\cdot),\cdot).$ (92)
Since $X_{\phi_{Y}}$ is the Hamiltonian part of Y and $\Lambda(dg,\cdot)$ is horizontal, we get $\Lambda(dg,\cdot)=\delta_{Y}$ and therefore equation (89) holds.
Conversely, let us consider $Y=X_{\phi_{Y}}+\Lambda(dg,\cdot)$. We want to show that $Y$ satisfies equations (53). A computation analogous to (91) yields
$L_{Y}\eta$ $=L_{X_{\phi_{Y}}}\eta+L_{\Lambda(dg,\cdot)}\eta$ (93)
$=-\mathcal{R}(\phi_{Y})\eta+L_{\left(\Lambda(dg,\cdot)-g\mathcal{R}\right)}\eta+L_{\left(g\mathcal{R}\right)}\eta$ (94)
$=-\mathcal{R}(\phi_{Y})\eta+L_{X_{g}}\eta+L_{g\mathcal{R}}\eta$ (95)
$=-\mathcal{R}(\phi_{Y})\eta-\mathcal{R}(g)\eta+gL_{\mathcal{R}}\eta+dg$ (96)
$=-\mathcal{R}(\phi_{Y})\eta-\mathcal{R}(g)\eta+dg$ (97)
$=-\mathcal{R}(\phi_{Y}+g)\eta+dg\,,$ (98)
thus showing that the first part of equations (53) holds with $a=-\mathcal{R}(\phi_{Y}+g)$. It remains to show the second condition in Theorem 1. An explicit computation gives
$Y(\mathcal{H})$ $=X_{\phi_{Y}}(\mathcal{H})+\Lambda(dg,d\mathcal{H})$ (99)
=$=\{\phi_{Y},\mathcal{H}\}_{\eta}-\mathcal{H}\mathcal{R}(\phi_{Y})+\{g,\mathcal{H}\}_{\eta}-\mathcal{H}\mathcal{R}(g)+g\mathcal{R}(\mathcal{H})$ (100)
=$=\{\phi_{Y}+g,\mathcal{H}\}_{\eta}-\mathcal{H}\mathcal{R}(\phi_{Y}+g)+g\mathcal{R}(\mathcal{H})$ (101)
$=a\mathcal{H}+g\mathcal{R}(\mathcal{H})+\{\phi_{Y}+g,\mathcal{H}\}_{\eta}.$ (102)
Since $\{\phi_{Y}+g,\mathcal{H}\}=0$ by hypothesis, we can conclude the proof. ∎
Corollary 5.
The invariant tensor densities associated with the Hamiltonian and horizontal components of a Cartan symmetry $Y$ are, respectively,
$-\phi_{Y}\eta^{-1}\quad\mbox{and}\quad(\eta\wedge dg)\eta^{-\frac{2}{n+1}}.$ (103)
The invariant tensor density for the Cartan symmetry is
$\mathfrak{K}=(\phi_{Y}+g)\ \eta^{-1}.$ (104)
Proof.
Proposition 10 implies that the Hamiltonian part of $Y$ is equivalent to the tensor density $-\phi_{Y}\eta^{-1}$. By Lemma 1 and Proposition 18, the Horizontal part is equivalent to
$\left(\eta\wedge dg\right)\eta^{-\frac{2}{n+1}}.$ (105)
Finally, using once more Proposition 18, we obtain the invariant tensor density. ∎
5 Integrability of contact Hamiltonian systems and their decomposition
In the classical Hamiltonian setting, Noether’s theorem gives a profound link between symmetries and conserved quantity, ultimately linking integrability to the presence of a sufficient number of independent constants of motion. In the contact case this is less clear, and in the past year attempts have already been made to give a notion of contact integrability and relate it with symmetries or dissipated quantities, see e.g. [de_Le_n_2020, Jovanovi__2015, Grillo_2020] and the references therein. In this section, we want to revisit this discussion in view of the decomposition presented above.
Definition 10.
A complete contact integrable system is a triple $\left(M,\eta,\{f_{0},\ldots,f_{n}\}\right)$, where $M$ is a $12n+$ 1-dimensional contact manifold endowed by a contact form $\eta$, and $\{f_{0},\ldots,f_{n}\}$ is a set of functions such that
  • $\{f_{i},f_{j}\}_{\eta}=0$ for all $i,j\in\{0,\ldots,n\}$
  • and $X_{f_{0}},\ldots,X_{f_{n}}$ are linear independent almost everywhere.
Remark 11.
In the symplectic version of Definition 10, the requirement of independence of the functions $\{h_{i}\}_{i=1,\ldots,n}$ is written as
[…]
This can be reformulated by considering the natural volume form induced by the symplectic form $\omega$ as
$\iota_{V_{h_{1}},\ldots,V_{h_{n}}}\omega^{\wedge n}:=\omega^{\wedge n}(V_{h_{1}},\ldots,V_{h_{n}},\cdot,\cdots,\cdot)\neq 0,$ (107)
where we denoted $V_{h}$ the (symplectic) Hamiltonian vector field associated to $h$, to distinguish it from their contact counterparts. See Appendix B for a small review of symplectic mechanics.
Example 2.
On a $4$-dimensional symplectic manifold $(P,\omega)$, a complete integrable system is defined by two functions in involution:
$\iota_{V_{f_{1}},V_{f_{2}}}\omega^{\wedge 2}$ $=2\ \omega(V_{f_{1}},V_{f_{2}})\omega+2\ \iota_{V_{f_{1}}}\omega\wedge\iota_{V_{f_{2}}}\omega$ (108)
$=2\{f_{1},f_{2}\}_{PB}\omega+2\ df_{1}\wedge df_{2}$ (109)
$=2\ df_{1}\wedge df_{2},$ (110)
where by {,}$\{\cdot,\cdot\}_{PB}$ we refer to the Poisson bracket induced by the symplectic form $\omega$.
The construction presented above extends to the contact case, were we can express the linear independence condition by via the natural volume form (1) induced by the contact form $\eta$, i.e.,
$\iota_{X_{f_{0}},X_{f_{1}},\ldots,X_{f_{n}}}(\eta\wedge(d\eta)^{\wedge n})\neq 0.$ (111)
The kernel of the $n$-form (111) corresponds to the distribution generated by the Hamiltonian vector field identified by the integrable system.
Moreover, the linear independence of a set of non-commuting vector fields $\{Y_{j}\}_{j\leq J}$ can be checked by the analogous condition
$\iota_{Y_{0},Y_{1},\ldots,Y_{J}}(\eta\wedge(d\eta)^{\wedge n})\neq 0.$ (112)
Lemma 4.
Let $(M,\eta)$ be a $(2n+1)$-dimensional exact contact manifold, and assume there is $m\leq n$ such that $\{f_{0},\ldots,f_{m}\}$ is a set of functions in involution whose contact Hamiltonian vector fields are linearly independent. Then
$\iota_{X_{f_{0}},X_{f_{1}},\ldots,X_{f_{n}}}\left(\eta\wedge(d\eta)^{\wedge n}\right)=$ (113)
$\quad=\frac{n!}{(n-m)!}\left(-1\right)^{\frac{1}{2}(m+1)(m+2)}\bigg{[}\sum_{j=0}^{m}\left(-1\right)^{jm}f_{j}\bigwedge_{r=1}^{m}df_{j+r\pmod{m+1}}\bigg{]}\wedge d\eta^{\wedge(n-m)}$ (114)
$\qquad+\frac{n!}{(n-m-1)!}\left(-1\right)^{\frac{1}{2}m(m+1)}\bigwedge_{j=0}^{m}df_{j}\wedge\eta\wedge d\eta^{\wedge(n-m-1)}.$ (115)
Proof.
We will prove the statement by induction. Let us consider $m=0$, so the set of functions consists only of $f_{0}$. The contraction of $X_{f_{0}}$ on the volume form is:
$\iota_{X_{f_{0}}}(\eta\wedge d\eta^{\wedge n})$ $=-f_{0}d\eta^{\wedge n}+n\ \left(df_{0}-\mathcal{R}(f_{0})\eta\right)\wedge\eta\wedge d\eta^{\wedge(n-1)}$ (116)
$=-f_{0}d\eta^{\wedge n}+n\ df_{0}\wedge\eta\wedge d\eta^{\wedge(n-1)}$ (117)
$=\left(-f_{0}d\eta+n\ df_{0}\wedge\eta\right)d\eta^{\wedge(n-1)}.$ (118)
Consider now $m=k-1$, the term would be
$\iota_{X_{f_{0}},X_{f_{1}},\ldots,X_{f_{k-1}}}\left(\eta\wedge(d\eta)^{\wedge n}\right)=$ (119)
$\quad=\left[\binom{n}{k-1}(k-1)!\right]\left(-1\right)^{\frac{1}{2}k(k+1)}\bigg{[}\sum_{j=0}^{k-1}\left(-1\right)^{j(k-1)}f_{j}\bigwedge_{r=1}^{k-1}df_{j+r\pmod{k}}\bigg{]}\wedge d\eta^{\wedge(n-k+1)}$ (120)
$\qquad+\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}k\left(k-1\right)}\bigwedge_{j=0}^{k-1}df_{j}\wedge\eta\wedge d\eta^{\wedge(n-k)}.$ (121)
First observe that
$\bigg{[}\sum_{j=0}^{k-1}\left(-1\right)^{j(k-1)}f_{j}\bigwedge_{r=1}^{k-1}df_{j+r\pmod{k}}\bigg{]}\left(X_{f_{k}}\right)$ (122)
$\quad=\bigg{[}\sum_{j=0}^{k-1}\left(-1\right)^{j(k-1)}\sum_{l=1}^{k-1}(-1)^{l}f_{j}X_{f_{k}}(f_{j+l\pmod{k+1}})\bigwedge_{\begin{subarray}{c}r=1\\ r\neq l\end{subarray}}^{k-1}df_{j+r\pmod{k}}\bigg{]}$ (123)
$\quad=\mathcal{R}(f_{k})\bigg{[}\sum_{j=0}^{k-1}\left(-1\right)^{j(k-1)}\sum_{l=1}^{k-1}(-1)^{l+1}f_{j}f_{j+l\pmod{k+1}}\bigwedge_{\begin{subarray}{c}r=1\\ r\neq l\end{subarray}}^{k-1}df_{j+r\pmod{k}}\bigg{]}=0,$ (124)
where between the second and the third line we exploited the vanishing of the Jacobi bracket between each pair of functions $\{f_{i}\}$ and equation (32). Thus, the contraction of the first expression in equation (119) with $X_{f_{k}}$ is
$\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}k\left(k+1\right)}\left(df_{k}-\mathcal{R}(f_{k})\eta\right)\wedge\bigg{[}\sum_{j=0}^{k-1}\left(-1\right)^{j(k-1)}f_{j}\bigwedge_{r=1}^{k-1}df_{j+r\pmod{k}}\bigg{]}\wedge d\eta^{\wedge(n-k)}.$ (125)
By contracting the second expression in equation (119) with the Hamiltonian vector field $X_{f_{k}}$, we obtain
$\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}k\left(k-1\right)}\left(\bigwedge_{j=0}^{k-1}df_{j}\wedge\eta\wedge d\eta^{\wedge n-k}\right)\left(X_{f_{k}}\right)$ (126)
$\quad=\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}k\left(k-1\right)}\bigg{(}\sum_{l=0}^{k-1}(-1)^{l+1}f_{l}\mathcal{R}(f_{k})\bigwedge_{\begin{subarray}{c}r=0\\ r\neq l\end{subarray}}^{k-1}df_{r}\wedge\eta\wedge d\eta^{\wedge(n-k)}$ (127)
$\qquad\qquad+(-1)^{k+1}f_{k}\bigwedge_{j=0}^{k-1}df_{j}\wedge d\eta^{\wedge(n-k)}$ (128)
$\qquad\qquad+(n-k)(df_{k})\bigwedge_{j=0}^{k-1}df_{j}\wedge\eta\wedge d\eta^{\wedge(n-k-1)}\bigg{)},$ (129)
in which we have used again the involutivity of the functions and equation (32). Combining the two results we obtain
$\iota_{X_{f_{0}},X_{f_{1}},\ldots,X_{f_{k}}}\left(\eta\wedge(d\eta)^{\wedge n}\right)=$ (130)
$\quad=\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}k\left(k+1\right)}\left(df_{k}\right)\wedge\bigg{[}\sum_{j=0}^{k-1}\left(-1\right)^{j(k-1)}f_{j}\bigwedge_{r=1}^{k-1}df_{j+r\pmod{k}}\bigg{]}\wedge d\eta^{\wedge(n-k)}$ (131)
$\qquad+\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}k\left(k-1\right)}\bigg{(}(-)^{k+1}f_{k}\bigwedge_{j=0}^{k-1}df_{j}\wedge d\eta^{\wedge(n-k)}$ (132)
$\qquad\qquad+(n-k)(df_{k})\bigwedge_{j=0}^{k-1}df_{j}\wedge\eta\wedge d\eta^{\wedge(n-k-1)}\bigg{)},$ (133)
after simplifying the two terms involving $\mathcal{R}(f_{k})$. Rearranging the first two terms, we obtain
$\iota_{X_{f_{0}},X_{f_{1}},\ldots,X_{f_{k}}}\left(\eta\wedge(d\eta)^{\wedge n}\right)=$ (134)
$\quad=\left[\binom{n}{k}k!\right]\left(-1\right)^{\frac{1}{2}(k+1)\left(k+2\right)}\bigg{[}\sum_{j=0}^{k}\left(-1\right)^{jk}f_{j}\bigwedge_{r=1}^{k}df_{j+r\pmod{k+1}}\bigg{]}\wedge d\eta^{\wedge(n-k)}$ (135)
$\qquad+\left[\binom{n}{k+1}(k+1)!\right]\left(-1\right)^{\frac{1}{2}k\left(k+1\right)}\bigwedge_{j=0}^{k}df_{j}\wedge\eta\wedge d\eta^{\wedge(n-k-1)}.$ (136)
Applying the identity $\binom{n}{m}m!=\frac{n!}{(n-m)!}$ then concludes the proof. ∎
Proposition 19.
Let $(M,\eta,\{f_{0},f_{1},\cdots,f_{n}\})$ be a complete integrable contact Hamiltonian system on a $(2n+1)$-dimensional contact manifold. Then
$\iota_{X_{f_{0}},X_{f_{1}},\ldots,X_{f_{n}}}\left(\eta\wedge(d\eta)^{\wedge n}\right)=n!(-1)^{\frac{1}{2}(n+1)(n+2)}\bigg{[}\sum_{j=0}^{n}\left(-1\right)^{jn}f_{j}\bigwedge_{r=1}^{n}df_{j+r\pmod{n+1}}\bigg{]}$ (137)
Proof.
Fix $m=n$, apply Lemma 4 and observe that the second expression in equation (113) vanishes. ∎
Remark 12.
In a three-dimensional contact manifold, a complete integrable contact Hamiltonian system is specified by $(M,\eta,\{f_{0},f_{1}\})$. It has a constant of motion, when it can be defined, $\kappa=f_{0}/f_{1}$. The $1$-form defined by equation (111) in this case reads
$\iota_{X_{f_{0}},X_{f_{1}}}\left(\eta\wedge d\eta\right)=-f_{0}df_{1}+f_{1}df_{0}+\{f_{0},f_{1}\}_{\eta}\ \eta=-f_{0}df_{1}+f_{1}df_{0}.$ (138)
This can be rewritten as
$\iota_{X_{f_{0}},X_{f_{1}}}\left(\eta\wedge d\eta\right)=-\frac{1}{f_{0}^{2}}\,d\kappa.$ (139)
Note that we use the involutivity only in the last identity in (138). The first identity holds for any pairs of Hamiltonian vector fields.
Integrability requires $f_{0}$ and $f_{1}$ to be linearly independent. However, note that in the submanifold $f_{0}=f_{1}=0$, if it exists, the two Hamiltonian vector fields $X_{f_{0}}$ and $X_{f_{1}}$ are always linearly dependent.
This means that a complete integrable contact Hamiltonian system cannot admit 0-levelsets. However, outside the 00-levelset, a system can be locally completely integrable. In this case, 00-levelsets identify equilibrium regions that could still be reached asymptotically. This can be described more generally in the sense of Sussmann integrability [Sussmann_1973].
Proposition 20.
Let $Y$ be a scaling symmetry of the Hamiltonian system $(M,\eta,\mathcal{H})$ and assume that there exists a contact Hamiltonian function $f$ in involution with $\mathcal{H}$, i.e., $\{f,\mathcal{H}\}_{\eta}=0$. Then
$\psi=\left\{\phi_{Y},f\right\}_{\eta}$ (140)
is a function in involution with $\mathcal{H}$.
Moreover, the scaling symmetry $Y$ is linearly independent from $X_{\mathcal{H}}$ and $X_{f}$ if
$\mathcal{H}\left(\psi+f\lambda\right)d\eta\wedge d\eta^{\wedge(n-1)}$ (141)
$\quad+(n-1)\bigg{[}(\mathcal{H}df-fd\mathcal{H})\wedge d\phi_{Y}+(\psi-\lambda\mathcal{H})df\wedge\eta+\phi_{Y}d\mathcal{H}\wedge df\bigg{]}\wedge d\eta^{\wedge(n-2)}$ (142)
$\quad+(n-1)(n-2)\;d\phi_{Y}\wedge d\mathcal{H}\wedge df\wedge\eta\wedge d\eta^{\wedge(n-3)}\neq 0$ (143)
Proof.
The first identity, (140), can be proved via the Jacobi identity,
$[Y,[X_{f},X_{\mathcal{H}}]]+[X_{\mathcal{H}},[X_{f},Y]]+[X_{f},[Y,X_{\mathcal{H}}]]=0.$ (144)
Applying the definition of scaling symmetry, this reduces to
$[X_{\mathcal{H}},[X_{f},Y]]=0.$ (145)
Then, it follows from Corollary 2 that $X_{\phi_{Y}}$ is itself a scaling symmetry and $\psi=\{\phi_{Y},f\}_{\eta}$ is in involution with $\mathcal{H}$.
Proposition 13 shows that $\psi=\{\phi_{Y},f\}_{\eta}$ is in involution with $\mathcal{H}$. Indeed,
00 0 $=\eta([X_{\mathcal{H}},[X_{f},Y]])$ (146)
$=L_{X_{\mathcal{H}}}(\eta([X_{f},Y]))-\iota_{[X_{f},Y]}L_{X_{\mathcal{H}}}\eta=$ (147)
$=L_{X_{\mathcal{H}}}\{f,\phi_{Y}\}+\mathcal{R}(\mathcal{H})\iota_{[X_{f},Y]}\eta=$ (148)
={,$=\{\mathcal{H},\{f,\phi_{Y}\}\}.$ (149)
We are left to prove the second part of the statement. Lemma 4 where $m=2$ and for the set of functions $\{\mathcal{H},f\}$ implies
$\iota_{X_{\mathcal{H}},X_{f}}\left(\eta\wedge d\eta^{\wedge n}\right)=-n\left[\left(\mathcal{H}df-fd\mathcal{H}\right)\wedge d\eta+(n-1)d\mathcal{H}\wedge df\wedge\eta\right]\wedge d\eta^{\wedge(n-2)}.$ (150)
Contracting equation (150) with $X_{\phi_{Y}}$, and using equations (140) and (60) we obtain:
$\iota_{X_{\mathcal{H}},X_{f}}\left(\eta\wedge d\eta^{\wedge n}\right)$ $=n\mathcal{H}\left(\psi+f\lambda\right)d\eta^{\wedge(n-1)}$ (151)
$\quad+(n-1)\bigg{[}(\mathcal{H}df-fd\mathcal{H})\wedge d\phi_{Y}$ (152)
$\quad\quad+(\psi-\lambda\mathcal{H})df\wedge\eta+\phi_{Y}d\mathcal{H}\wedge df\bigg{]}\wedge d\eta^{\wedge n-2}$ (153)
$\quad+(n-1)(n-2)d\phi_{Y}\wedge d\mathcal{H}\wedge df\wedge\eta\wedge d\eta^{\wedge(n-3)};$ (154)
since $n\neq 0$ (otherwise the contact manifold would be $1$-dimensional) and $d\eta\neq 0$, this concludes the proof. ∎
A similar result holds in the more general case of dynamical similarities. In this case, however, it does not directly hold for function in involution but for constants of motion.
Proposition 21.
Let $Y\in\Gamma(TM)$ be a dynamical similarity of the contact Hamiltonian system $(M,\eta,\mathcal{H})$. If there exists a function $f\in C^{\infty}(M)$ that is a constant of motion of the Hamiltonian, i.e., $X_{\mathcal{H}}f=0$, then $\psi=Y\left({f}\right)$ is also a constant of motion for $\mathcal{H}$.
Proof.
By definition $[Y,X_{\mathcal{H}}](f)=gX_{\mathcal{H}}(f)=0$, that is, $X_{\mathcal{H}}\left(Y(f)\right)=0$.  ∎
Acknowledgments
The authors would like to thank Alessandro Bravetti and Angel Alejandro García-Chung for the useful discussions. The research of MS and FZ was supported by the NWO project 613.009.10.
[…]
Appendix A Tensor Densities
Here we are going to sketch the definition of tensor densities and some of their properties. For a more detailed discussion, we refer to [Fuks1986] and [projective, Chapter 6.3].
Let $M$ be a smooth manifold of dimension $m$. The space of tensor densities of degree $\lambda\in\mathbb{R}$, denoted by $F_{\lambda}$, is the space of sections of the bundle $\left(\bigwedge^{m}T^{*}M\right)^{\lambda}$, i.e. $\mathcal{F}_{\lambda}=\Gamma(\left(\bigwedge^{m}T^{*}M\right)^{\lambda})$. Locally in coordinates, any such generalized tensor $\phi$ can be expressed at a point $x\in M$ as
$\phi(x)=f(x)\ (dx_{1}\wedge\cdots\wedge dx_{n})^{\lambda},$ (155)
for some $\lambda\in\mathbb{R}$ and $f\in C^{\infty}(M)$. An diffeomorphism $g\in\mathrm{Diff}(M)$ acts on tensor densities $\phi$ by
$g^{-1}\phi=\left(Dg\right)^{\lambda}\ \left(\phi\circ g\right),$ (156)
where $Dg$ is the Jacobian of the transformation $g$.
If $\zeta\in\Omega^{m}(M)$ is a volume form over $M$, a tensor density of degree $\lambda$ can be written in local coordinates as
$\phi(x)=f(x)\ \zeta^{\lambda}.$ (157)
If we consider a contact manifold $(M,\eta)$ of dimension $n$, it is interesting to restrict the attention to its group of contactomorphisms [projective, Chapter 7.5]. Since a contactomorphism transforms the natural volume form $\eta\wedge(d\eta)^{\wedge n}$ as (9) and the contact form $\eta$ as (8), we can view the contact form as a tensor density of degree $\lambda=\frac{1}{n+1}$:
$\eta=\left(\eta\wedge(d\eta)^{\wedge n}\right)^{\frac{1}{n+1}}.$ (158)
We can then express any tensor density $\phi$ of degree $\lambda\in\mathbb{R}$ on a contact manifold as
$\phi(x)=f(x)\left(\eta\wedge(d\eta)^{\wedge n}\right)^{\lambda}=f(x)\eta^{\lambda(n+1)}.$ (159)
Appendix B Symplectic Hamitonian systems
Let $N$ be an even dimensional manifold and $\omega$ a non degenerate closed two-form, we call the couple $(N,\omega)$ a symplectic manifold [Arnold2009, sympgeo] and $\omega$ a symplectic form.
A real function $h\in C^{1}(N)$ is called an Hamiltonian function, usually this function encodes the energy of a mechanical physical system. Given an Hamiltonian function $h$, the symplectic forms allows to induce a vector field $V_{h}$ by
$\omega(V_{h},\cdot)=dh,$ (160)
called (symplectic) Hamiltonian vector field. We call the triple $(N,\omega,h)$ a symplectic Hamiltonian system to distinguish it from the contact Hamiltonian systems studied in the paper. Since $\omega$ is skew-symmetric, one has
$\omega(V_{h},V_{h})=V_{h}(h)=0.$ (161)
Observing that $\dot{H}=\omega(V_{h},V_{h})$, intended as the time derivative along the flow of $V_{h}$, one can rephrase the identity above as conservation of energy.
Definition 11.
A vector field $\Xi$ on a symplectic manifold $(N,\omega)$ is a Liouville vector field if it satisfies
$L_{\Xi}\omega=\omega.$ (162)